1.05 g of NaOH pellet is added to 45 mL of distilled water and the temperature rises from 25.1 oC to 35.8 oC. Determine the heat per mole of NaOH generated in this process. Assume that the density of water is 1 g/mL and the specific heat capacity of the solution is 4.184 J/goC
Answer #1

moles of NaOH = 1.05 / 40 = 0.02625

dT = 35.8 -25.1 = 10.7 oC

total mass of solution = 1.05 + 45 = 46.05 g

Q   = m Cp dT

Q = 46.05 x 4.184 x 10.7

Q = 2062 J

Q = 2.062 kJ

heat per mole NaOH = - 2.062 / 0.02625

                                    = - 78.5 kJ / mol